OpenStudy (anonymous):

What is the y-intercept of the function y = 1,500(2.15)x? x 2.15 1,500 1

jimthompson5910 (jim_thompson5910):

Same idea. Plug in x = 0 and evaluate.

OpenStudy (anonymous):

where would i plug it in thats where im confused

jimthompson5910 (jim_thompson5910):

Wherever you see an x y = 1,500(2.15)^x y = 1,500(2.15)^0

OpenStudy (anonymous):

I got 1? is that correct?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

I did it again and i got 1,500 is that correct?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

its 2,500

OpenStudy (precal):

Jim_Thompson5910 is correct